# MMedAgent-RL: Optimizing Multi-Agent Collaboration for Multimodal Medical Reasoning

### v2

@@ Appendix A Evaluated Models
We evaluate a series of state-of-the-art LVLMs and Multi-agent. The single-agent models include LLaVA [30], Yi-VL-34B [61], Qwen-VL [3], LLaVA-Med [24], MedFlamingo [36], RadFM [54] and GPT-4o [39]. The multi-agent frameworks include prior collaborative systems such as MedAgents [48] and MDAgents [20], as well as our proposed MMedAgent-RL framework that introduces reinforcement learning for adaptive multi-agent reasoning.
@@
  • GPT-4o [39] is OpenAI’s latest multimodal large model that supports text, image, and audio inputs. It exhibits strong generalization across vision-language benchmarks and serves both as a single-agent baseline and as a specialist in our multi-agent settings.
@@
  • Med-Flamingo [36] is a multimodal few-shot learner designed for the medical domain. Built upon OpenFlamingo, it is further pre-trained on biomedical image-text data from scientific literature. It enables few-shot medical visual question answering with minimal supervision.
@@
  • RadFM [54] is a domain-specific foundation model tailored for radiology. It leverages large-scale radiology reports and domain-adaptive learning to improve zero-shot and few-shot performance on radiographic image understanding.
@@
  • LLaVA-Med [24] extends LLaVA to the biomedical domain by fine-tuning with medical image-instruction pairs. It enhances medical reasoning and answer generation with limited supervision using domain-specific visual-textual alignments.
@@
  • Qwen2.5-VL [3] is a versatile vision-language model developed by Alibaba. It supports high-quality OCR, multi-turn dialogue, and reasoning over complex multimodal inputs. It is used both as a strong single-agent baseline and as the foundation of agents in our proposed framework.
@@
  • Yi-VL-34B [61] is a large-scale multimodal model from 01.AI. With 34 billion parameters, it offers high-capacity visual understanding and serves as a powerful open-source baseline across medical and general VQA tasks.
@@
  • LLaVA [31, 30] are general-purpose vision-language models trained via visual instruction tuning. Evaluated in several sizes (7B, 13B, 34B), they serve as strong single-agent baselines in both in-domain and out-of-domain medical benchmarks.
Appendix B Evaluated Datasets
@@
We employ three established medical vision-language datasets: VQA-RAD [22], SLAKE [29], and PathVQA [15]. Furthermore, to evaluate out-of-distribution performance, we incorporate the health and medicine subset of MMMU [62] along with OmniMedVQA [18].
@@
  • VQA-RAD [22] is a manually constructed dataset containing 315 radiology images with 3,515 question-answer pairs. The images are distributed across head, chest, and abdomen regions, and include both open-ended and binary "yes/no" questions. Each image is associated with multiple clinically relevant questions generated by medical professionals. The dataset aims to facilitate the development of visual question answering systems for the medical domain.
@@
  • SLAKE [29] is a semantically-labeled knowledge-enhanced dataset featuring 642 radiology images and over 14,000 question-answer pairs. It offers comprehensive annotations including masks for semantic segmentation and bounding boxes for object detection. SLAKE is bilingual (English and Chinese) and covers 12 diseases and 39 organs across various body parts. The dataset also incorporates a medical knowledge graph with 5,232 medical knowledge triplets to support knowledge-based reasoning.
@@
  • PathVQA [15] is a pathology-focused dataset containing 32,799 open-ended questions from 4,998 pathology images. The dataset was created using a semi-automated pipeline to extract images and captions from pathology textbooks and generate question-answer pairs using natural language processing. PathVQA aims to support the development of AI systems capable of answering clinical questions about pathology images, with each question manually checked for correctness.
@@
  • MMMU [62] (Health & Medicine subset) is part of the Massive Multi-discipline Multimodal Understanding benchmark. This subset contains approximately 1,752 test questions across five disciplines: Basic Medical Science, Clinical Medicine, Diagnostics and Laboratory Medicine, Pharmacy, and Public Health. The questions require college-level subject knowledge and deliberate reasoning, challenging models to perform expert-level perception and reasoning tasks.
@@
  • OmniMedVQA [18] is a comprehensive medical VQA benchmark collected from 73 different medical datasets, featuring images across 12 different modalities and covering more than 20 distinct anatomical regions. All images are sourced from authentic medical scenarios, ensuring alignment with real-world applications. The benchmark provides a diverse evaluation platform for testing the capabilities of large vision-language models in medical image understanding and reasoning.
Appendix C Overview of the Baselines
@@
We evaluate MMedAgent-RL against two main multi-agent baselines, MedAgents [48] and MDAgents [20]. These baselines represent state-of-the-art approaches in medical visual question answering.
@@
  • MedAgents [48] establishes a zero-shot multi-agent collaboration framework that simulates real-world clinical workflows. The framework encompasses five critical steps: gathering domain experts, proposing individual analyses, summarizing analyses into a report, iterating over discussions until consensus is reached, and making a final decision. Different agents are assigned specific medical roles and collaborate to solve complex medical reasoning tasks. The framework relies on pre-trained large language models without additional fine-tuning, enabling natural dialogue-based interactions between agents. MedAgents demonstrates how specialized medical knowledge from different domains can be integrated through structured agent collaboration, providing a strong baseline for multi-agent medical reasoning.
@@
  • MDAgents [20] advances multi-agent medical systems by introducing adaptive collaboration mechanisms. Unlike fixed collaboration patterns, MDAgents dynamically selects the most appropriate agent configuration and communication structure based on the specific medical task. This framework allows for more flexible interactions between general practitioners and specialist agents, optimizing the collaboration pattern for different types of medical queries. MDAgents incorporates mechanisms to resolve conflicts between different agent opinions and adapts the consultation workflow to match the complexity of the medical case, resulting in more robust decision-making across diverse medical scenarios.
Appendix D Experimental Setup
D.1 Data Statistics
@@
- Easy 8,321 498 1,284 6,539 / /
- Medium 1,409 160 114 1,135 / /
- Hard 2,626 281 275 2,070 / /
@@
D.2 Hyperparameter Settings
We use Qwen2.5-VL [3] as the base model. We design the prompt template using the format employed in MM-EUREKA [35], clearly specifying the required output structure, which includes using <think> and <answer> tags to separately contain the reasoning process and the final answer, with the two being separated. The detailed prompt is shown in Table 4. For training hyperparameters, the rollout batch size and training batch size are both set to 128, with 8 rollouts generated for each sample. The sampling temperature is set to 1.0 to encourage response diversity, and optimization is done with a learning rate of $1\times 10^{-6}$. Additionally, for the three stages of curriculum reinforcement learning, the KL divergence coefficients are set to $1\times 10^{-3}$, $4\times 10^{-3}$, and $1\times 10^{-2}$ respectively to stabilize training. For the number of specialists, we set $e=3$. For the baseline implementation, i.e., MedAgents [48] and MDAgents [20], we use Qwen2.5-VL as the agent for decision making to ensure a fair comparison between multi-agent baselines and MMedAgent-RL. For model training of 3B and 7B, we use DeepSpeed ZeRO2 and ZeRO3 respectively [43]. For the training framework, we adopt a multimodal RL framework based on OpenRLHF [17]. All training is conducted on 8 NVIDIA Tesla A100 80GB GPUs.
D.3 Prompt
@@
Appendix E Additional Results
E.1 Ablation Analysis
E.1.1 Performance of Triage Doctor
Table 5: The performance of triage doctor.
Model VQA-RAD SLAKE PathVQA
Qwen2.5-VL-3B 95.62 92.16 77.53
Qwen2.5-VL-7B 96.21 94.41 80.58
MMedAgent-RL (3B) 99.96 99.94 98.56
MMedAgent-RL (7B) 99.98 99.96 99.12
E.1.2 KL Divergence Coefficient
Figure 7: Ablation of KL divergence coefficient.
E.2 Scaling Up the Model Parameters to 7B
Comparison with Baselines in In-Distribution Datasets. Table 6 also presents the performance of 7B-scale models across the same in-distribution medical VQA benchmarks. Overall, 7B models exhibit consistent improvements over their 3B counterparts. Notably, Qwen2.5-VL-7B outperforms Qwen2.5-VL-3B across all in-domain datasets, achieving an average score of 62.3% compared to 60.4%. This trend is further amplified in collaborative multi-agent setups. Specifically, MMedAgent-RL (7B) achieves a new state-of-the-art in in-domain settings with a 71.2% average, surpassing its 3B version by 3.3 percentage points and outperforming all other collaborative baselines. These results underscore the scalability of our collaborative learning strategy and reinforce the importance of model capacity in complex medical reasoning tasks.
Performance in Out-of-Distribution Datasets. In OOD settings, larger models also demonstrate improved generalization. As shown in Table 6, MMedAgent-RL (7B) achieves an average score of 60.7%, representing a 5.8% gain over the best-performing single-agent 7B model (Qwen2.5-VL-7B, 54.9%). Compared to its 3B counterpart, it improves by 4.3%, demonstrating that scaling not only benefits in-domain tasks but also enhances robustness to domain shifts. The performance margins over MedAgents and MDAgents are particularly notable—MMedAgent-RL (7B) outperforms these baselines by 8.2% and 5.6%, respectively—highlighting the crucial role of reinforcement learning in optimizing collaborative decision making in medical VQA under distributional shifts.
Table 6: The results of the medical VQA benchmark.
@@ E.2 Scaling Up the Model Parameters to 7B
Model VQA-RAD SLAKE PathVQA Avg. OmniMedVQA MMMU Avg.

### v3

@@ Appendix C Evaluated Models
We evaluate a series of state-of-the-art LVLMs and Multi-agent. The single-agent models include LLaVA (Liu et al., 2024a), Yi-VL-34B (Young et al., 2024), Qwen-VL (Bai et al., 2025), LLaVA-Med (Li et al., 2023), MedFlamingo (Moor et al., 2023), RadFM (Wu et al., 2023), HuatuoGPT-Vision (Chen et al., 2024a) and GPT-4o (OpenAI, 2024). The multi-agent frameworks include prior collaborative systems such as MedAgents (Tang et al., 2024), MDAgents (Kim et al., 2024) and AFlow (Zhang et al., 2024), as well as our proposed MMedAgent-RL framework that introduces reinforcement learning for adaptive multi-agent reasoning.
@@
  • GPT-4o (OpenAI, 2024) is OpenAI’s latest multimodal large model that supports text, image, and audio inputs. It exhibits strong generalization across vision-language benchmarks and serves both as a single-agent baseline and as a specialist in our multi-agent settings.
@@
  • Med-Flamingo (Moor et al., 2023) is a multimodal few-shot learner designed for the medical domain. Built upon OpenFlamingo, it is further pre-trained on biomedical image-text data from scientific literature. It enables few-shot medical visual question answering with minimal supervision.
@@
  • RadFM (Wu et al., 2023) is a domain-specific foundation model tailored for radiology. It leverages large-scale radiology reports and domain-adaptive learning to improve zero-shot and few-shot performance on radiographic image understanding.
@@
  • LLaVA-Med (Li et al., 2023) extends LLaVA to the biomedical domain by fine-tuning with medical image-instruction pairs. It enhances medical reasoning and answer generation with limited supervision using domain-specific visual-textual alignments.
@@
  • Qwen2.5-VL (Bai et al., 2025) is a versatile vision-language model developed by Alibaba. It supports high-quality OCR, multi-turn dialogue, and reasoning over complex multimodal inputs. It is used both as a strong single-agent baseline and as the foundation of agents in our proposed framework.
@@
  • Yi-VL-34B (Young et al., 2024) is a large-scale multimodal model from 01.AI. With 34 billion parameters, it offers high-capacity visual understanding and serves as a powerful open-source baseline across medical and general VQA tasks.
@@
  • LLaVA (Liu et al., 2024b; a) are general-purpose vision-language models trained via visual instruction tuning. Evaluated in several sizes (7B, 13B, 34B), they serve as strong single-agent baselines in both in-domain and out-of-domain medical benchmarks.
  • HuatuoGPT-Vision-7B (Chen et al., 2024a) is a medical multimodal large language model (MLLM) trained on the curated PubMedVision dataset. This dataset was created by using GPT-4V to denoise and reformat 1.3 million image-text pairs from PubMed, significantly improving data quality. As a result, HuatuoGPT-Vision demonstrates superior performance on medical multimodal benchmarks compared to other open-source models.
Appendix D Evaluated Datasets
@@
We employ three established medical vision-language datasets: VQA-RAD (Lau et al., 2018), SLAKE (Liu et al., 2021), and PathVQA (He et al., 2020). Furthermore, to evaluate out-of-distribution performance, we incorporate the health and medicine subset of MMMU (Yue et al., 2024) along with OmniMedVQA (Hu et al., 2024b).
@@
  • VQA-RAD (Lau et al., 2018) is a manually constructed dataset containing 315 radiology images with 3,515 question-answer pairs. The images are distributed across head, chest, and abdomen regions, and include both open-ended and binary "yes/no" questions. Each image is associated with multiple clinically relevant questions generated by medical professionals. The dataset aims to facilitate the development of visual question answering systems for the medical domain.
@@
  • SLAKE (Liu et al., 2021) is a semantically-labeled knowledge-enhanced dataset featuring 642 radiology images and over 14,000 question-answer pairs. It offers comprehensive annotations including masks for semantic segmentation and bounding boxes for object detection. SLAKE is bilingual (English and Chinese) and covers 12 diseases and 39 organs across various body parts. The dataset also incorporates a medical knowledge graph with 5,232 medical knowledge triplets to support knowledge-based reasoning.
@@
  • PathVQA (He et al., 2020) is a pathology-focused dataset containing 32,799 open-ended questions from 4,998 pathology images. The dataset was created using a semi-automated pipeline to extract images and captions from pathology textbooks and generate question-answer pairs using natural language processing. PathVQA aims to support the development of AI systems capable of answering clinical questions about pathology images, with each question manually checked for correctness.
@@
  • MMMU (Yue et al., 2024) (Health & Medicine subset) is part of the Massive Multi-discipline Multimodal Understanding benchmark. This subset contains approximately 1,752 test questions across five disciplines: Basic Medical Science, Clinical Medicine, Diagnostics and Laboratory Medicine, Pharmacy, and Public Health. The questions require college-level subject knowledge and deliberate reasoning, challenging models to perform expert-level perception and reasoning tasks.
@@
  • OmniMedVQA (Hu et al., 2024b) is a comprehensive medical VQA benchmark collected from 73 different medical datasets, featuring images across 12 different modalities and covering more than 20 distinct anatomical regions. All images are sourced from authentic medical scenarios, ensuring alignment with real-world applications. The benchmark provides a diverse evaluation platform for testing the capabilities of large vision-language models in medical image understanding and reasoning.
Appendix E Overview of the Baselines
@@
We evaluate MMedAgent-RL against two main multi-agent baselines, MedAgents (Tang et al., 2024), MDAgents (Kim et al., 2024) and AFlow (Zhang et al., 2024). These baselines represent state-of-the-art approaches in medical visual question answering.
@@
  • MedAgents (Tang et al., 2024) establishes a zero-shot multi-agent collaboration framework that simulates real-world clinical workflows. The framework encompasses five critical steps: gathering domain experts, proposing individual analyses, summarizing analyses into a report, iterating over discussions until consensus is reached, and making a final decision. Different agents are assigned specific medical roles and collaborate to solve complex medical reasoning tasks. The framework relies on pre-trained large language models without additional fine-tuning, enabling natural dialogue-based interactions between agents. MedAgents demonstrates how specialized medical knowledge from different domains can be integrated through structured agent collaboration, providing a strong baseline for multi-agent medical reasoning.
@@
  • MDAgents (Kim et al., 2024) advances multi-agent medical systems by introducing adaptive collaboration mechanisms. Unlike fixed collaboration patterns, MDAgents dynamically selects the most appropriate agent configuration and communication structure based on the specific medical task. This framework allows for more flexible interactions between general practitioners and specialist agents, optimizing the collaboration pattern for different types of medical queries. MDAgents incorporates mechanisms to resolve conflicts between different agent opinions and adapts the consultation workflow to match the complexity of the medical case, resulting in more robust decision-making across diverse medical scenarios.
  • AFlow (Zhang et al., 2024) is a framework designed to automatically generate and optimize complex problem-solving workflows for LLMs. Instead of relying on a single inference pass, these workflows enhance performance through structured procedures. We evaluate three distinct strategies as strong baselines: Self-Consistency Ensemble, which runs an agent multiple times and selects the most frequent answer to improve reliability; Multi-Agent Debate, which uses several agents to collaboratively propose, critique, and refine solutions; and Self-Refine, which employs a feedback loop for a single agent to iteratively critique and improve its own output.
Appendix F Experimental Setup
F.1 Data Statistics
@@
   - Easy 8,321 498 1,284 6,539 / /
   - Medium 1,409 160 114 1,135 / /
   - Hard 2,626 281 275 2,070 / /
@@
F.2 Hyperparameter Settings
We use Qwen2.5-VL (Bai et al., 2025) as the base model. We design the prompt template using the format employed in MM-EUREKA (Meng et al., 2025), clearly specifying the required output structure, which includes using <think> and <answer> tags to separately contain the reasoning process and the final answer, with the two being separated. The detailed prompt is shown in Table 4. For training hyperparameters, the rollout batch size and training batch size are both set to 128, with 8 rollouts generated for each sample. The sampling temperature is set to 1.0 to encourage response diversity, and optimization is done with a learning rate of $1\times 10^{-6}$. Additionally, for the three stages of curriculum reinforcement learning, the KL divergence coefficients are set to $1\times 10^{-3}$, $4\times 10^{-3}$, and $1\times 10^{-2}$ respectively to stabilize training. The dynamic entropy coefficient $\gamma_{s}$ is set to $0.03$ for hard ($s=0$), $0.005$ for medium ($0<s<1$), and $0.0001$ for easy ($s=1$) samples, respectively, to adapt the level of exploration based on curriculum difficulty. For the number of specialists, we set $e=3$. For the baseline implementation, i.e., MedAgents (Tang et al., 2024), MDAgents (Kim et al., 2024) and AFlow (Zhang et al., 2024), we use Qwen2.5-VL as the agent for decision making to ensure a fair comparison between multi-agent baselines and MMedAgent-RL. For the training framework, we adopt a multimodal RL framework based on OpenRLHF (Hu et al., 2024a). For the inference, we adopt the vLLM framework (Kwon et al., 2023). All training is conducted on 8 NVIDIA Tesla A100 80GB GPUs.
F.3 Prompt
@@
F.4 Dataset Splitting and Difficulty Stratification
The partition of training, validation, and test sets strictly follows the official dataset guidelines. The difficulty grading is required exclusively for the training phase to implement curriculum learning. Specifically, we classify samples into difficulty levels based on the consistency and accuracy of the specialists’ responses. We emphasize that the inference process and the calculation of evaluation scores do not use this difficulty grading at all. The model generates responses without access to any difficulty labels or ground truth. The difficulty stratification on the test set is applied solely for the analytical breakdown shown in Figure 4. It serves only to visualize and analyze model performance across different complexity levels post-generation, and has no influence on the inference process itself. The difficulty grading for the test dataset follows the same as training data.
Appendix G Additional Results
G.1 Comparison with Supervised Fine-Tuning (SFT)
We compare our method with Supervised Fine-Tuning (SFT) methods as a baseline. As shown in Table 5, our proposed method, MMedAgent-RL, demonstrates a significant performance advantage over SFT methods across all the medical VQA benchmarks. The superiority of our approach is even more pronounced in the more challenging out-of-distribution datasets. This highlights our model’s enhanced robustness and generalization capabilities. Overall, MMedAgent-RL consistently sets a new state-of-the-art, with our base model achieving a 73.3% average on in-domain tasks and 72.6% on out-of-distribution tasks, already surpassing the SFT method. The results clearly indicate that our multi-agent, reinforcement learning-based approach is more effective than traditional SFT techniques for complex medical VQA tasks.
Table 5: The comparison with SFT on several medical VQA benchmarks.
@@ G.1 Comparison with Supervised Fine-Tuning (SFT)
Model VQA-RAD SLAKE PathVQA Avg. OmniMedVQA MMMU-Med Avg.